OpenStudy (anonymous):

Which of the following represents x^2-16/x^2-4x in simplified form?

OpenStudy (anonymous):

A. 1-4/x B. 1+4/x C.x+4/x D.x-4/x

OpenStudy (anonymous):

@Callisto @jazy

OpenStudy (callisto):

\[ \frac{x^2-16}{x^2-4x} = \frac{(x+4)(x-4)}{x(x-4)} =\frac{x+4}{x}\]

OpenStudy (anonymous):

that makes more sense and thank you!

OpenStudy (callisto):

Welcome :)